Could somebody please help me with: -2(1-3x)+3(5x+2)=109
Answer by lwsshak3(11628)

Could somebody please help me with: -2(1-3x)+3(5x+2)=109
**
-2(1-3x)+3(5x+2)=109
clear parenthesis and distribute
-2+6x+15x+6=109
combine like terms
21x+4=109
add -4 to both sdes
21x+4-4=109-4
21x=105
divide both sides by 21
x=105/21
x=5